If 24.15 ml of sodium hydroxide were required in the titration of a 4.9651 g sample of dried reagent potassium biphthalate (KHC8H4O4), what would be the normality of sodium hydroxide?


1
Expert's answer
2022-04-21T11:22:05-0400

NaOH + KHC8H4O4 = NaKC8H4O4 + H2O

n = m / M

M (KHC8H4O4) = 204.22 g/mol

n (KHC8H4O4) = 4.9651 / 204.22 = 0.024 mol

n (NaOH) = n (KHC8H4O4) = 0.024 mol

CM (NaOH) x V (NaOH) = n (KHC8H4O4)

CM (NaOH) = n (KHC8H4O4) / V (NaOH) = 0.024 / 0.02415 = 1 M

When doing normal/molar conversions, remember when the valence is one, normality and molarity are the same. The normality of the 1 M solution of NaOH is 1 N.